DETERMINE THE NUMBER OF SETS OF ALL POSSIBLE RANDOM SAMPLES THAT CAN BE DRAWN FROM THE GIVEN POPULATION BY USING THE FORMULA, NCn.N = 5, n = 3.

1
Expert's answer
2022-05-13T03:28:35-0400

The number of possible samples which can be selected without replacement is

"_NC_n=\cfrac{N! } {n! \cdot(N-n)! };"

"_{5}C_3=\cfrac{5! } {3! \cdot(5-3)! }=\cfrac{4\cdot5}{2}=10."